# arXiv:2602.23440v1[cs.CL] 26 Feb 2026

### v1

@@ Abstract
Training large language models to reason with search engines via reinforcement learning is hindered by a fundamental credit assignment problem: existing methods such as Search-R1 provide only a sparse outcome reward after an entire multi-step trajectory, making it infeasible to attribute success or failure to individual reasoning and retrieval decisions. Process-reward methods like StepSearch alleviate this by introducing step-level supervision, but rely on heuristic rewards such as TF-IDF overlap with gold documents, and still sample $k$ complete trajectories per example, retaining high gradient variance. We propose Slate, a framework built on two complementary ideas: (1) truncated step-level sampling, which generates $k$ trajectories that share a common prefix and differ only at the next step, and (2) dense LLM-as-judge rewards, which replace heuristic scoring with a capable LLM evaluator that assesses the quality of each reasoning step, search query, and answer, providing richer and more reliable supervision. We theoretically prove that under the same dense reward structure, truncated sampling reduces the variance of advantage estimates by up to a factor of $T$ compared to full-trajectory sampling for $T$-step trajectories, yielding lower-variance, better-targeted policy gradients. Experiments on seven QA benchmarks confirm that Slate consistently outperforms both sparse-reward and process-reward baselines, with the largest gains on harder multi-hop tasks and smaller models.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation (Hendrycks et al., 2020; Clark et al., 2018). Despite these achievements, LLMs often struggle with complex reasoning tasks (Wei et al., 2022) and lack access to up-to-date external knowledge (Jin et al., 2024). Integrating search engines into the LLM reasoning loop, where the model interleaves its own chain-of-thought reasoning with external retrieval calls, has emerged as a promising paradigm for knowledge-intensive question answering (Yao et al., 2023; Trivedi et al., 2022a).
Reinforcement learning (RL) provides a natural framework for optimizing such search-augmented reasoning systems. Search-R1 (Jin et al., 2025) pioneered the use of RL (PPO and GRPO) to train LLMs that autonomously invoke search engines during multi-turn reasoning, using outcome-based exact-match rewards. While effective, this sparse reward design suffers from a fundamental limitation: the credit assignment problem. When the model receives a single binary signal after completing a multi-step trajectory, potentially involving several rounds of thinking, query generation, and information processing, it cannot determine which intermediate steps contributed positively or negatively to the outcome. Recent work has attempted to address this limitation through process-level supervision. StepSearch (Wang et al., 2025) augments PPO with step-wise rewards based on information gain and redundancy penalties, achieving notable improvements on multi-hop QA benchmarks. However, StepSearch still relies on sampling $k$ complete trajectories per training example, which is computationally expensive and retains variance from all steps in the gradient estimates for any individual step.
@@
  • Dense LLM-as-Judge Rewards: We replace the sparse EM outcome reward with dense, step-level rewards produced by an LLM evaluator. The judge scores each reasoning step (thinking quality), each search query (query quality), and the final answer (correctness) on a discrete scale $\{-1,0,+1\}$, providing rich supervision at every decision point.
We provide formal theoretical analysis showing that truncated step-level sampling yields advantage estimates with provably lower variance than standard full-trajectory sampling, with up to a $T$-fold reduction for $T$-step trajectories, resulting in lower-variance policy gradients. Combined with the dense reward signal from the LLM judge, this enables faster convergence, better credit assignment, and improved final performance. Our contributions are the following:
@@
    We propose truncated step-level sampling for GRPO, an exploration strategy that generates $k$ alternative continuations only at the current decision point rather than $k$ complete trajectories. We formally prove this achieves a $T$-fold advantage variance reduction over full-trajectory sampling, yielding lower-variance policy gradients.
@@
    We introduce an LLM-as-judge reward system that provides dense, discrete step-level supervision for reasoning quality, search query quality, and answer correctness, replacing the sparse outcome reward.
@@
    We demonstrate that Slate significantly outperforms both existing sparse-reward methods (Search-R1) and process-reward methods (StepSearch) on seven QA benchmarks.
@@ 2 Related Work
Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) integrates external knowledge into LLM generation by retrieving relevant passages based on the input query. While effective for simple lookups, standard RAG struggles with multi-hop reasoning where iterative retrieval is needed (Yang et al., 2018; Trivedi et al., 2022a). Alternative approaches treat the search engine as a tool (Yao et al., 2023; Schick et al., 2023), enabling the LLM to decide when and what to search. However, prompting-based tool-use methods rely on in-context examples and fail to generalize robustly, while supervised fine-tuning approaches require expensive annotated trajectories (Schick et al., 2023; Asai et al., 2024).
@@
Reinforcement learning has emerged as a potent paradigm for enhancing LLM reasoning, as demonstrated by OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025). Policy gradient methods including PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) have been widely adopted for training reasoning models. GRPO is particularly attractive as it eliminates the need for a separate critic model by using group-relative advantages computed from multiple sampled responses. However, applying these methods to search-augmented scenarios introduces unique challenges around reward design, credit assignment, and integration of retrieved content.
@@
Search-R1 (Jin et al., 2025) introduced the paradigm of RL training with interleaved search engine calls, using outcome-based EM rewards and retrieved token loss masking. Follow-up work includes R1-Searcher (Song et al., 2025), ReSearch (Chen et al., 2025), and ZeroSearch (Sun et al., 2025), all of which rely on sparse global rewards. StepSearch (Wang et al., 2025) addresses the reward sparsity problem by introducing step-wise rewards with information gain and redundancy penalties, but still samples complete trajectories and relies on access to ground-truth intermediate documents for computing information gain. Our work differs from all prior methods by combining truncated step-level sampling with model-based dense rewards from an LLM judge (eliminating the need for ground-truth intermediate annotations).
@@ 3.3 Dense LLM-as-Judge Rewards
The query reward $r_{\text{query}}(q_{t},s_{t},\tau_{<t})\in\{-1,0,+1\}$ evaluates the search query $q_{t}$ conditioned on the retrieved documents, the preceding reasoning and the trajectory context. The judge assesses five criteria: relevance, specificity, searchability, alignment, and novelty. A score of +1+1 indicates a specific, well-formed query with clear keywords; 0 indicates some specificity but generic phrasing; and 1-1 indicates an irrelevant or poorly formed query.
@@ 4 Theoretical Analysis
Since the policy gradient is a linear function of these advantages, lower advantage variance directly translates to lower-variance gradient estimates, enabling faster convergence and better final solutions (see Appendix A.6 for a detailed discussion).
Remark 1 (Credit Assignment).
@@ 5.4 Main Results
(1) Slate consistently achieves the best performance across all benchmarks. On the 7B model, Slate obtains an average EM of 0.461, representing a 3.0% absolute (7.0% relative) improvement over Search-R1 (0.431) and outperforming the best prior results on every individual dataset. On the 3B model, the improvement over Search-R1 is even more substantial (0.396 vs. 0.303, a 30.7% relative improvement), demonstrating that smaller models benefit more from the dense step-level supervision.
@@
(3) Modest but consistent gains on general QA. On the general QA benchmarks, Slate outperforms Search-R1 by 1.3–1.9% absolute EM, with smaller gains expected for shorter trajectories where credit assignment is less challenging.
(4) Smaller models benefit proportionally more from dense supervision. On the 3B model (Table 3), gains over Search-R1 are dramatically larger on multi-hop benchmarks (e.g., +16.7% on Musique, +27.3% on Bamboogle), suggesting smaller models benefit most from explicit step-level supervision.
@@
5.5 Analysis
5.5.1 Ablation Study
Table 2: Ablation study on Qwen2.5-7B-Base (Exact Match).
@@ 5.5.1 Ablation Study
To understand the contribution of each component, we conduct an ablation study on Qwen2.5-7B-Base (Table 2). We evaluate four variants: (a) Slate without truncated sampling (using $k$ full trajectories with LLM-judge rewards), (b) Slate without LLM-judge rewards (using truncated sampling with sparse EM reward), (c) Slate with EM reward only at the final step (no thinking/query rewards), and (d) the full Slate system. The results reveal that both components contribute meaningfully, with their benefits most pronounced on harder datasets. Removing truncated sampling (variant a) reduces average EM by 1.1%, with the largest drops on the hardest benchmarks (Musique -1.1%, Bamboogle -1.3%), confirming that step-level variance reduction from shared prefixes improves optimization quality especially for complex multi-hop reasoning. Removing LLM-judge rewards (variant b) causes a larger drop of 2.4% average EM, demonstrating that dense step-level rewards provide a stronger training signal than sparse EM. The impact is again largest on the harder datasets (Musique -2.9%, Bamboogle -3.7%). Variant (c), which uses truncated sampling but only provides EM reward at the final step, performs close to Search-R1 (0.368 vs. 0.361 avg.), showing that truncated sampling alone without dense rewards provides limited benefit, it is the combination of both that yields the full improvement.
Figure 2: Training dynamics comparison on Qwen2.5-7B-Base. Slate converges faster and achieves a higher, more stable reward compared to Search-R1/GRPO and StepSearch/StePPO.
5.5.2 Training Dynamics
We compare the training reward curves of Slate against Search-R1 (GRPO) and StepSearch (StePPO) in Figure 2. Slate exhibits three notable properties: (1) Faster convergence: Slate reaches its peak training reward approximately 20% faster than StepSearch, attributable to the denser gradient signal from step-level rewards. (2) Higher reward ceiling: The final training reward of Slate is consistently higher than both baselines, reflecting the improved credit assignment from truncated sampling. (3) Greater stability: Unlike GRPO which can exhibit reward collapse, Slate maintains stable optimization throughout training due to the lower-variance advantage estimates predicted by Theorem 1.
5.5.3 Effect of Group Size $k$
We study the impact of $k\in\{1,3,5,7\}$ truncated samples per step (Appendix A.9). Performance improves steadily from k=1k{=}1 to k=5k{=}5 with diminishing returns at k=7k{=}7, consistent with the $1/k$ variance reduction from Eq. 4.
@@ 6 Conclusion
We presented Slate, a novel training method for search-augmented reasoning that introduces truncated step-level sampling and dense LLM-as-judge rewards. By sampling $k$ trajectory continuations that share a common prefix and differ only at the current decision step, our method achieves provably lower advantage variance than standard full-trajectory sampling, with up to a $T$-fold reduction for $T$-step trajectories, leading to lower-variance policy gradients. The dense LLM-as-judge rewards replace sparse binary outcome signals with rich, fine-grained step-level supervision that directly evaluates reasoning quality, query quality, and answer correctness on a ternary scale.
Our theoretical analysis provides formal guarantees on variance reduction, and our empirical results demonstrate significant improvements over both sparse-reward methods (Search-R1) and process-reward methods (StepSearch) across seven question answering benchmarks. Slate establishes a new state of the art for RL-based search-augmented reasoning, demonstrating that the combination of step-level exploration and dense model-based rewards is a powerful paradigm for training LLMs that reason effectively with external knowledge sources.
@@ A.1 Algorithm
20:  $\theta\leftarrow\theta+\alpha\nabla_{\theta}\Big[\frac{1}{k}\sum_{j=1}^{k}\mathcal{J}_{t}^{(j)}(\theta)-\beta\,\mathbb{D}_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\Big]$
@@
31:end for
Method NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle Avg.
Direct 0.152 0.434 0.178 0.145 0.249 0.018 0.030 0.172
CoT 0.168 0.451 0.196 0.163 0.257 0.032 0.072 0.191
RAG 0.312 0.504 0.296 0.251 0.221 0.051 0.076 0.244
IRCoT 0.286 0.478 0.268 0.171 0.196 0.059 0.245 0.243
Search-o1 0.267 0.465 0.254 0.240 0.207 0.045 0.316 0.256
SFT 0.301 0.491 0.286 0.191 0.248 0.039 0.112 0.238
R1 (no search) 0.258 0.467 0.243 0.191 0.277 0.051 0.232 0.246
ZeroSearch 0.348 0.531 0.365 0.260 0.234 0.056 0.096 0.270
ReSearch 0.339 0.518 0.342 0.261 0.228 0.074 0.184 0.278
Search-R1 0.406 0.587 0.435 0.284 0.273 0.049 0.088 0.303
StepSearch 0.329 0.339 0.181 0.328
Slate (Ours) 0.425 0.603 0.451 0.351 0.368 0.216 0.361 0.396

### v2

# Truncated Step-Level Sampling with Process Rewards for Retrieval-Augmented Reasoning

@@ Abstract
Training large language models to reason with search engines via reinforcement learning is hindered by a fundamental credit assignment problem: existing methods such as Search-R1 provide only a sparse outcome reward after an entire multi-step trajectory, making it infeasible to attribute success or failure to individual reasoning and retrieval decisions. Process-reward methods like StepSearch alleviate this by introducing step-level supervision, but rely on heuristic rewards such as TF-IDF overlap with gold documents, and still sample $k$ complete trajectories per example, retaining high gradient variance. We propose Slate, a framework built on two complementary ideas: (1) truncated step-level sampling, which generates $k$ trajectories that share a common prefix and differ only at the next step, isolating variation to a single decision point; and (2) dense, decomposed LLM-as-judge rewards, which score each reasoning step, search query, and answer on a ternary scale with separate quality dimensions, providing richer supervision than binary outcome signals or undifferentiated step-level judgments. We theoretically prove that under the same dense reward structure, truncated sampling reduces the variance of advantage estimates by up to a factor of $T$ compared to full-trajectory sampling for $T$-step trajectories, yielding lower-variance and better-targeted policy gradients. Experiments on seven QA benchmarks confirm that Slate consistently outperforms both sparse-reward and process-reward baselines, with the largest gains on harder multi-hop tasks and smaller models.
@@ 1 Introduction
Large language models have demonstrated remarkable capabilities in natural language understanding and generation (Hendrycks et al., 2020; Clark et al., 2018). Despite these achievements, LLMs often struggle with complex reasoning tasks (Wei et al., 2022) and lack access to up-to-date external knowledge (Jin et al., 2024). Integrating search engines into the LLM reasoning loop, where the model interleaves its own chain-of-thought reasoning with external retrieval calls, has emerged as a promising paradigm for knowledge-intensive question answering (Yao et al., 2023; Trivedi et al., 2022a).
Reinforcement learning provides a natural framework for optimizing such systems. Search-R1 (Jin et al., 2025) pioneered RL training for LLMs that invoke search engines during multi-turn reasoning, using outcome-based exact-match rewards. While effective, this sparse reward suffers from the credit assignment problem: a single binary signal after a multi-step trajectory cannot attribute success or failure to individual steps. Process-level supervision methods, StepSearch (Wang et al., 2025) with heuristic step rewards, and SWiRL (Goldie et al., 2025) with LLM-judge binary step rewards, improve over outcome-only supervision, but both still sample complete trajectories independently, meaning step-level advantages conflate variation from the current action with variation across different prefix histories, leaving the credit assignment and gradient variance problems only partially resolved.
@@
  • Dense LLM-as-Judge Rewards: We replace the sparse EM outcome reward with dense, step-level rewards produced by an LLM evaluator. The judge scores each reasoning step (thinking quality), each search query (query quality), and the final answer (correctness) on a discrete scale, providing rich supervision at every decision point.
Our contributions are the following:
@@
    We propose truncated step-level sampling for GRPO, an exploration strategy that generates $k$ alternative continuations from a shared prefix at the current decision point rather than $k$ complete trajectories. We formally prove this achieves a $T$-fold advantage variance reduction over full-trajectory sampling, yielding lower-variance policy gradients. This principled exploration mechanism is, to our knowledge, the first to provide formal guarantees on variance reduction for step-level RL in search-augmented reasoning.
@@
    We introduce a decomposed LLM-as-judge reward system with separate ternary scores for reasoning quality, query quality, and answer correctness, along with an early-termination bonus. Unlike prior binary step-level judgments (Goldie et al., 2025), this structured design provides richer supervision that disentangles the distinct skills required for effective retrieval-augmented reasoning.
@@
    We demonstrate that Slate significantly outperforms both existing sparse-reward methods (Search-R1) and process-reward methods (StepSearch) on seven QA benchmarks, and that our ablation confirms truncated sampling provides gains beyond what step-level rewards alone achieve.
@@ 2 Related Work
Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) integrates external knowledge into LLM generation but struggles with multi-hop reasoning requiring iterative retrieval (Yang et al., 2018; Trivedi et al., 2022a). Tool-use approaches (Yao et al., 2023; Schick et al., 2023) let the LLM decide when and what to search, but prompting-based methods fail to generalize robustly and supervised fine-tuning requires expensive annotated trajectories (Schick et al., 2023; Asai et al., 2024).
@@
Reinforcement learning has become a key paradigm for LLM reasoning (Jaech et al., 2024; Guo et al., 2025), with policy gradient methods such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) widely adopted. GRPO is particularly attractive as it eliminates the need for a separate critic by using group-relative advantages. However, applying these methods to search-augmented scenarios introduces challenges around reward design, credit assignment, and integration of retrieved content.
@@
Search-R1 (Jin et al., 2025) introduced the paradigm of RL training with interleaved search engine calls, using outcome-based EM rewards and retrieved token loss masking. Follow-up work includes R1-Searcher (Song et al., 2025), ReSearch (Chen et al., 2025), and ZeroSearch (Sun et al., 2025), all of which rely on sparse global rewards. StepSearch (Wang et al., 2025) addresses the reward sparsity problem by introducing step-wise rewards with information gain and redundancy penalties, but still samples complete trajectories and relies on access to ground-truth intermediate documents for computing information gain. SWiRL (Goldie et al., 2025) explored using an LLM judge to provide step-level binary rewards for multi-step reasoning with tool use. However, SWiRL operates offline, generating complete trajectories, segmenting them into sub-trajectories, and training on this static corpus, so step-level advantages still conflate variation across different prefix histories with variation at the current step. It also uses undifferentiated binary step judgments and provides no formal variance analysis. Our work differs along three axes: (1) truncated step-level sampling generates $k$ continuations from an identical shared prefix, isolating variation to exactly one decision point with provable variance guarantees (Theorem 1); (2) online GRPO optimization avoids the distribution mismatch of offline approaches; and (3) our decomposed ternary reward design provides richer supervision than heuristic step rewards (Wang et al., 2025) or binary LLM judgments (Goldie et al., 2025), without requiring ground-truth intermediate annotations.
@@ 3.3 Dense LLM-as-Judge Rewards
The query reward $r_{\text{query}}(q_{t},s_{t},\tau_{<t})\in\{-1,0,+1\}$ evaluates the search query $q_{t}$ conditioned on the preceding reasoning and the trajectory context. The judge assesses five criteria: relevance, specificity, searchability, alignment, and novelty. A score of +1+1 indicates a specific, well-formed query with clear keywords; 0 indicates some specificity but generic phrasing; and 1-1 indicates an irrelevant or poorly formed query. By design, the query reward evaluates the formulation quality of the query before observing retrieval results; this decoupling ensures the reward reflects intrinsic query properties (keyword specificity, search-engine compatibility) rather than fluctuating with retrieval nondeterminism or index coverage, while the downstream effect of retrieval quality is captured through subsequent steps and the final answer reward.
@@ 4 Theoretical Analysis
Note: The two estimators target different quantities, the trajectory-level advantage $\hat{A}_{i}$ estimates the deviation of the full return $R(\tau)$, while the step-level advantage $\hat{A}_{t}^{(j)}$ estimates the deviation of the single-step reward $r_{t}$. The comparison below therefore characterizes a bias-variance trade-off: the step-level estimator achieves lower variance at the cost of discarding future-reward information (see Remark 1 for when this substitution is justified).
@@
Since the policy gradient is a linear function of these advantages, lower advantage variance directly translates to lower-variance gradient estimates, enabling faster convergence and better final solutions (see Appendix A.7 for a detailed discussion). We further discuss the bias-variance trade-off of estimator substitution and the credit assignment benefits of dense rewards in Appendix A.4.
@@ 5.1 Main Results
(1) Slate consistently performs best across all benchmarks. On the 7B model, Slate obtains an average EM of 0.461, representing a 3.0% absolute (7.0% relative) improvement over Search-R1 (0.431) and outperforming the best prior results on every individual dataset. On the 3B model, the improvement over Search-R1 is even more substantial (0.396 vs. 0.303, a 30.7% relative improvement), demonstrating that smaller models benefit more from the dense step-level supervision.
@@
(3) Gains on general QA and out-of-domain generalization. On the general QA benchmarks, Slate outperforms Search-R1 by 1.3–1.9% absolute EM. Notably, five of our seven evaluation benchmarks are out-of-domain (trained only on NQ+HotpotQA), and the largest gains occur on these unseen multi-hop tasks, demonstrating that step-level supervision teaches transferable reasoning skills rather than dataset-specific shortcuts.
(4) Smaller models benefit more from dense supervision. On the 3B model (Table 3), gains over Search-R1 are dramatically larger on multi-hop benchmarks (e.g., +16.7% on Musique, +27.3% on Bamboogle), suggesting smaller models benefit most from explicit step-level supervision.
@@
Ablation Study
Table 2: Ablation study on Qwen2.5-7B-Base (Exact Match). Variant (a) mirrors the SWiRL (Goldie et al., 2025) approach of applying step-level LLM-judge rewards with full-trajectory sampling (adapted to our online setting).
@@
Table 2 ablates each component of Slate on Qwen2.5-7B-Base. Variant (a) is particularly informative: it applies dense LLM-judge rewards with full-trajectory sampling, closely mirroring the SWiRL (Goldie et al., 2025) approach (adapted to our online setting). While this already improves over Search-R1 (+2.9% avg.), the full Slate with truncated sampling achieves a further +1.1% gain, with the largest improvements on the hardest benchmarks (Musique ++1.1%, Bamboogle ++1.3%). This demonstrates that truncated sampling provides gains above and beyond step-level rewards alone, confirming that the exploration strategy, not just the reward signal, is critical, consistent with Theorem 1. Removing LLM-judge rewards (variant b) causes a larger drop of 2.4% average EM, with the biggest impact on harder datasets (Musique -2.9%, Bamboogle -3.7%). Variant (c), with only EM reward at the final step, performs close to Search-R1 (0.368 vs. 0.361). Note that variant (c) still receives a non-trivial training signal at intermediate steps: among the $k$ candidates sampled at each step, some may produce an <answer> (receiving the EM reward) while others produce a <search> (receiving zero reward), creating a nonzero advantage that trains the model on when to stop searching. However, the small margin over Search-R1 confirms that it is the synergy of both components that yields the full improvement.
Figure 2: Training dynamics comparison on Qwen2.5-7B-Base.
Training Dynamics
We compare the training reward curves of Slate against Search-R1 (GRPO) and StepSearch (StePPO) in Figure 2. Slate exhibits three notable properties: (1) Faster convergence: Slate reaches its peak training reward approximately 20% faster than StepSearch, attributable to the denser gradient signal from step-level rewards. (2) Higher reward ceiling: The final training reward of Slate is consistently higher than both baselines, reflecting the improved credit assignment from truncated sampling. (3) Greater stability: Unlike GRPO which can exhibit reward collapse, Slate maintains stable optimization throughout training due to the lower-variance advantage estimates predicted by Theorem 1. We also study the effect of the number of truncated samples $k$ per step in Appendix A.12. Performance improves steadily from $k=1$ to $k=5$ with diminishing returns at $k=7$, consistent with the $1/k$ variance reduction predicted by Theorem 1.
Why Process Rewards Succeed in Search.
In math, step-level rewards as RL training signals generally underperform outcome-based rewards (Uesato et al., 2022; Lightman et al., 2023). Search-augmented reasoning differs in several key ways: steps are externally grounded by retrieval results (eliminating capability mismatch), rewards are evaluative rather than predictive, and trajectories are short ($T\leq 4$) and modular. Truncated sampling further addresses prefix confounding by isolating variation to a single action (Theorem 1). We provide a detailed discussion in Appendix A.11.
@@ 6 Conclusion
We presented Slate, a training method for search-augmented reasoning whose key insight is that how step-level optimization is performed matters as much as what reward signal is used. By sampling $k$ continuations from a shared prefix, our method isolates variation to a single action with provable variance guarantees (up to $T$-fold reduction), a formal contribution absent from prior step-level reward approaches (Goldie et al., 2025; Wang et al., 2025). Combined with decomposed ternary rewards, Slate significantly outperforms both sparse-reward and process-reward baselines across seven benchmarks, with ablations confirming that truncated sampling yields gains above and beyond step-level rewards alone.
@@ A.1 Algorithm
20:  Accumulate gradient: $\Delta\theta\mathrel{+}=\nabla_{\theta}\Big[\frac{1}{k}\sum_{j=1}^{k}\mathcal{J}_{t}^{(j)}(\theta)-\beta\,\mathbb{D}_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}]\Big]$
@@
[…]
32:end for
Note: Gradients are accumulated across all steps within a trajectory and applied in a single batched update after the trajectory completes. In practice, updates are further batched across multiple examples in the training batch.